Solve the following system of linear equation addition method
8x+3y=-5
3x+3y=10

Answer by 20800563(2)

3x+3y=10 let multiply this equation with -1 and then obtain that -3x-3y=-10 then let add this one:
8x+3y=-5
-3x-3y=-10
+__________
5x=-15 and x=-3 so when we put x for 8x+3y=-5 we have -24+3y=-5 and then we have 3y=19 so y=19/3 and we get solution (-3,19/3)